OpenStudy (anonymous):

A bottle of salon shampoo costs twenty-five dollars for twenty ounces. What is the unit price of the shampoo? eighty cents per bottle one dollar and twenty-five cents per ounce zero point eight zero ounces per dollar one point two five ounces per bottle

OpenStudy (whpalmer4):

Unit price is the cost per unit. If you pay $25 and get 20 oz, how much do you pay for each ounce? Some of the answers are clearly ridiculous choices. One of them is a number you might get if you set the problem up incorrectly.

OpenStudy (anonymous):

$1.25

OpenStudy (anonymous):

right @whpalmer4 ?

OpenStudy (whpalmer4):

$1.25 per ounce, not just $1.25.

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

can you help me on more questions?

OpenStudy (whpalmer4):

It's a fraction: \[\frac{$25}{20 oz} = \frac{25}{20} $/oz = $1.25/oz\] Sure, ask away...

OpenStudy (anonymous):

Estimate which benchmark percentage is closest to nine-seventeenths. twenty-five percent fifty percent seventy-five percent one hundred percent

OpenStudy (whpalmer4):

\[\frac{9}{17} =\] Let's look at our benchmarks: \[25\% = \frac{25}{100}\]\[50\%=\frac{50}{100}\]\[75\%=\frac{75}{100}\]\[100\%=\frac{100}{100}\]

OpenStudy (anonymous):

25%

OpenStudy (whpalmer4):

What do you have to multiply 17 by to get 100? We're estimating, so let's make the arithmetic easier by rounding up to 20.

OpenStudy (whpalmer4):

20 times what gives you 100?

OpenStudy (anonymous):

20 times 5

OpenStudy (whpalmer4):

Okay, good! So, if we multiply the top and the bottom of the fraction by the same number, we don't change its value, right? We're just multiplying by 1, written differently. What do we get if we multiply \[\frac{9}{17}\]by 5 on both top and bottom?

OpenStudy (whpalmer4):

Make it easier: multiply 9/20 top and bottom...

OpenStudy (whpalmer4):

that gives us \[\frac{45}{100}\] right? Which of our benchmark percentages is that closest to?

OpenStudy (anonymous):

50%

OpenStudy (whpalmer4):

Right! As it turns out, 9/17 = 0.5294... so that's the answer we want. Here's a different way we could have gotten there, given that we had choices and simply had to select one. 9/17 let's try 25%: 0.25*17 = 4.25 Nope, that's too small, we want something around 9 let's try 75% 0.75*17 = 12.25 nope, that's too big, we want something around 9 let's try 50% 0.50*17 = 8.5 that's pretty close! 9/17 must be about 50%.

OpenStudy (anonymous):

OK I need these last 11 real fast cause I'm going to have to go in like 5 minutes so please help really fast. :)

OpenStudy (whpalmer4):

Another way to roll: all of those benchmark percentages are fractions: 25 % = 1/4 50 % = 2/4 = 1/2 75 % = 3/4 100 % = 4/4 How could we change 9/17 to make it close to one of those fractions? Well, how about adding 1 to the denominator, giving us 9/18? That reduces easily to 1/2, so 50% is our benchmark estimate again.

OpenStudy (anonymous):

Estimate by rounding to the nearest half. four and one-seventh plus five and five-eighths minus two and one-ninth Answer choices: 6 six and one-half 7 seven and one-half

OpenStudy (whpalmer4):

4 1/7 - nearest half is 4, right? 5 5/8 - nearest half is 5 1/2, right? 2 1/9 - nearest half is 2, right? 4 + 5 1/2 - 2 = ?

OpenStudy (anonymous):

7

OpenStudy (anonymous):

7 and a half!

OpenStudy (whpalmer4):

Yes! (to 7 1/2)

OpenStudy (anonymous):

Marlo receives a ten percent bonus on a salary of forty-five thousand dollars. What is his bonus? two hundred twenty-two dollars four hundred fifty dollars two thousand two hundred dollars four thousand five hundred dollars

OpenStudy (whpalmer4):

10% = 10/100. What is 10/100 * $45,000?

OpenStudy (anonymous):

1,000

OpenStudy (whpalmer4):

No. That isn't one of the answer choices, not even one of the wrong ones :-)

OpenStudy (anonymous):

2,200?

OpenStudy (whpalmer4):

What is 45,000 * 10?

OpenStudy (whpalmer4):

Multiply or divide by a power of 10 just changes the decimal point, the digits don't change...

OpenStudy (whpalmer4):

Come on. If you had $5 and I offered to take your $5 and give you 10 times as much money, how much would I have to give you?

OpenStudy (anonymous):

55

OpenStudy (anonymous):

450..

OpenStudy (whpalmer4):

Multiply by 10: write down your number, add another 0. 10*10 = 100 45*10 = 450 45,000*10 = 450,000 Divide by 100: write down your number, move the decimal point to the left 2 places. 1000/100 = 10.00 10000/100 = 100.00 450000/100 = 4500.00 (you misread my proposal slightly - I didn't offer to give you your money back, just 10 times as much as you had, so $50 was the right answer)

OpenStudy (whpalmer4):

Another way of thinking of 10% is just that it is 1/10 of the amount. So if he got 10% on $45,000, that's just $45,000/10 = $4,500.

OpenStudy (anonymous):

Robyn paid $18 for a 2-month movie club membership. What is the unit rate? $6 per month $9 per month $18 per month $20 per month

OpenStudy (whpalmer4):

$18 paid, 2 months. How much per month?

OpenStudy (whpalmer4):

(It cost me $19 to see a movie yesterday with my sweetie, not counting popcorn — I want to know where I sign up for this $18 for 2 months deal :-)

OpenStudy (anonymous):

9dollars per month

OpenStudy (whpalmer4):

Right!

OpenStudy (anonymous):

Which is the simplified form of 6 to 2? 1 to 4 one-third 3 to 1 four colon one

OpenStudy (anonymous):

I believe it's 1 over 3

OpenStudy (whpalmer4):

No, have to keep them in the same order...if the big one came first, it has to be first in the simplified form, too

OpenStudy (anonymous):

1 to 4?

OpenStudy (whpalmer4):

\[\frac{6}{2}\]Are there any numbers that you could factor out of both top and bottom?